OpenStudy (anonymous):

Complete pattern 1 n 10

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

... wha...?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Oops here is the pattern 1,2,4,7,11,--,--

OpenStudy (anonymous):

16,22

OpenStudy (anonymous):

you're adding 1, then 2, then 3, then 4

OpenStudy (anonymous):

you 1+1=2 2+2=4 4+3=7 7+4=11 11+5=16 16+6=22

OpenStudy (anonymous):

^ yessir